OpenStudy (anonymous):

What is the order of √5,-0.1, -5/3 , 0.7, √2 from least to greatest?

OpenStudy (igreen):

\(\sqrt 5 = 2.24\) \(\dfrac{-5}{~~3} = -1.67\) \(\sqrt 2 = 1.41\) So we have: \(2.24\) \(-0.10\) \(-1.67\) \(0.70\) \(1.41\) Arrange in least to greatest: \(-1.67, -0.10, 0.70, 1.41, 2.24 \rightarrow \dfrac{-5}{~~3}, -0.1, 0.7, \sqrt 2, \sqrt 5\)

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @KCpaige